in the right triangle, the ratio of the measures of the two acute angles is 4:1. what is the measure in degrees, of the larger acute angle?
Answer by checkley77(12844)

4N+1N=5N
5N=90
N=90/5
N=18 DEGREES.
4N=4*18=72 DEGREES ANS.
PROOF
18+72+90=180
180=180